OpenStudy (anonymous):

181.25 is what percent of 29?

OpenStudy (aravindg):

(X/100)*29=181.25 Solve for X

OpenStudy (anonymous):

I am confused

OpenStudy (aravindg):

Take a paper and solve this out. I gave you the method.

OpenStudy (tkhunny):

29 is 100% of 29 -- Notice that 29*1 = 29 58 is 200% of 29 -- Notice that 29*2 = 58 116 is 400% of 29 -- Notice that 29*4 = 116